# Characterizing rational homogeneous spaces via $\Csuperscript\C\C^{*}$-actions

### v1

@@ 5.1. The orbit graph
Recalling that a fixed point of ${\mathcal{M}}$ parametrizes either a fixed curve or a bridge we see that the fixed point components of the action on ${\mathcal{M}}$ are images via $p$ of the varieties ${\mathcal{U}}_{j}^{i}$, $(i=-,0,+)$. We will then set
@@
Let ${\mathcal{N}}$ be an irreducible component of ${\mathcal{M}}_{j}^{0}$. By Remark 2.9, ${\mathcal{N}}$ is an irreducible component of the normalization of $\mathop{\rm RatCurves}\nolimits(Y_{j})$, hence it is a family of rational curves in $Y_{j}$, and it is obviously unsplit. Moreover, since $T_{Y_{j}}$ is a direct summand of ${T_{X}}_{|Y_{j}}$ (see Section 2.1), it follows that $\nu^{*}T_{Y_{j}}$ is nef, for the normalization $\nu$ of any curve of the family ${\mathcal{N}}$, and we conclude that ${\mathcal{N}}$ is beautiful by Remark 2.9. ∎
@@
The proof of this statement will obtained by combining Theorem 5.3 together with Lemmata 5.8, 5.9, 5.10 below.
@@ 5.2. Geometry of the BB-cells
In this section we will discuss some special properties of the plus and minus BB-decompositions associated with an equalized ${\mathbb{C}}^{*}$-action on a smooth projective variety of Picard number one admitting a beautiful family of rational curves. We will work in the situation of Setup 5.1. We start by showing that in this situation such decompositions are stratifications, i.e. that the closure of a BB-cell is a union of BB-cells:
@@
It is clearly enough to prove only the first part of the statement, and to prove it only for $j=i+1$. In order to do so, we will show that a general orbit with sink in $Y_{i+1}$ is contained in $B_{i}^{-}$. Note that Theorem 5.7 implies that there exists an orbit with sink in $Y_{i+1}$ and source in $Y_{r}$, from which we may conclude that $B_{i+1}^{-}$ intersects the open subset $X^{+}(Y_{r})$ and so the general orbit with sink in $Y_{i+1}$ has source in $Y_{r}$. Let $\Gamma$ be the closure of such an orbit, and set ${\mathcal{U}}_{\Gamma}:=q^{-1}(\Gamma)$.
@@
Consider the ruled surface $S_{C}$ as in Construction 1. The limit curve $C_{-}$ is a bridge connecting $Y_{i}$ and $Y_{i+1}$, while the limit curve $C_{+}$ is either a curve contained in $Y_{r}$, or a bridge connecting $Y_{r-1}$ and $Y_{r}$ (if $Y_{r}$ is an isolated point). In all cases, the sink of the action on $S_{C}$ is a point mapped into $X$ to a point of $Y_{i}$, and the source is mapped to $Y_{r}$. Hence, a general orbit in $S_{C}$ is mapped to an orbit contained in $B^{-}(Y_{i})$, and therefore $q(S_{C})\subset B^{-}(Y_{i})$; since $\Gamma=q(\Gamma^{\prime})\subset q(S_{C})$ the proof is finished. ∎
@@
In the first case $\Gamma^{\prime}$ connects the source of $\Gamma_{-}$ with the source of $\Gamma_{+}$ and we are in case 0b) of Proposition 4.6, in the second $\Gamma^{\prime}$ connects the source of $\Gamma_{-}$ with a point of $\Gamma_{+}$, which is fixed, and we are in case 1) of Proposition 4.6. In both cases the cycle $\Gamma_{-}+\Gamma^{\prime}$ is a limit of closure of general orbits in $S_{\Gamma}$. These orbits are mapped to orbits with sink $y$ and source belonging to $Y_{r}$. ∎
@@
For every $i=0,\dots,r-1$ the intersection $B_{i}^{-}\cap B_{i+1}^{+}$ is proper. i.e., $\mathop{\rm codim}\nolimits(B_{i}^{-}\cap B_{i+1}^{+})=\mathop{\rm codim}\nolimits B_{i}^{-}+\mathop{\rm codim}\nolimits B_{i+1}^{+}=\nu^{+}_{i}+\nu^{-}_{i+1}$.
@@
Let us consider again the Chow quotient $\mathcal{C}\!X$ of $X$ by the ${\mathbb{C}}^{*}$-action (cf. [23, Section 4]), which is an $(n-1)$-dimensional projective variety.
Set $B:=B_{i}^{-}\cap B_{i+1}^{+}$. Since (by Theorem 5.3) there exists a bridge between $Y_{i}$ and $Y_{i+1}$, then $B$ intersects the open sets $X^{-}(Y_{i})\subset B_{i}^{-}$, $X^{+}(Y_{i+1})\subset B_{i+1}^{+}$, and we may claim that the ${\mathbb{C}}^{*}$-orbit of a general point in $B$ has its extremal fixed point components in $Y_{i},Y_{i+1}$. We then conclude that $B$ is the locus of the set of bridges between $Y_{i}$ and $Y_{i+1}$, that we denoted by $I$; note that $I={\mathcal{U}}_{i}^{+}={\mathcal{U}}_{i+1}^{-}$. In particular $B$ is irreducible and $\dim B=\dim I+1$.
We denote by $e_{i}:I\to Y_{i}$ (resp. $e_{i+1}:I\to Y_{i+1}$) the map assigning to every element of $I$ its sink (resp. its source). Let us denote by $V_{i}^{+}\subset{\mathbb{P}}(N(Y_{i})^{+})$, (resp. $V_{i+1}^{-}\subset{\mathbb{P}}(N(Y_{i+1})^{-})$) the open set parametrizing orbits linking $Y_{0}$ and $Y_{i}$ (resp. $Y_{i+1}$ and $Y_{r}$), which is nonempty by Theorem 5.7. Consider the fiber product:
@@
which parametrizes all the connected ${\mathbb{C}}^{*}$-invariant $1$-cycles of the form $C^{\prime}+C+C^{\prime\prime}$, where $C^{\prime}$ is the closure of an orbit linking $Y_{0}$, $Y_{i}$, $C\in I$ and $C^{\prime\prime}$ is the closure of an orbit linking $Y_{i+1}$, $Y_{r}$. The variety $Z$ is smooth, irreducible, by construction, of dimension:
@@
On the other hand we may consider in $\mathcal{C}\!X$ the subvariety parametrizing connected ${\mathbb{C}}^{*}$-invariant cycles of the form $C_{1}+C_{2}$ with $C_{1}$ linking $Y_{0}$ and $Y_{i}$, $C_{2}$ linking $Y_{i}$ and $Y_{r}$. It is isomorphic to an open subset of ${\mathbb{P}}(N(Y_{i})^{-})\times_{Y_{i}}{\mathbb{P}}(N(Y_{i})^{+})$. We denote its closure by $E_{i}\subset\mathcal{C}\!X$, and define analogously $E_{i+1}$. By construction $E_{i}$, $E_{i+1}$ are irreducible divisors in $\mathcal{C}\!X$.
@@
In order to prove that $Z\subset E_{i}\cap E_{i+1}$ we have to show that every cycle of $Z$ belongs to $E_{i}\cap E_{i+1}$. Such cycles are of the form $C^{\prime}+C+C^{\prime\prime}\in Z$, where $C^{\prime},C,C^{\prime\prime}$ are ${\mathbb{C}}^{*}$-invariant irreducible curves linking, respectively, $Y_{0}$ to $Y_{i}$, $Y_{i}$ to $Y_{i+1}$, and $Y_{i+1}$ to $Y_{r}$. It is then enough to notice that, by Lemma 5.12 $C+C^{\prime\prime}$ deforms algebraically to a ${\mathbb{C}}^{*}$-invariant curve linking $Y_{0}$ to $Y_{i+1}$ and, analogously, that $C^{\prime}+C$ deforms algebraically to a ${\mathbb{C}}^{*}$-invariant curve linking $Y_{i}$ to $Y_{r}$. ∎
An important outcome of the previous statement, that we will use later on, is that it allows us to relate the ranks $\nu^{\pm}_{i}$ to the dimensions of the extremal fixed point components of the action on the varieties ${\mathcal{M}}_{y}$, $y\in X^{{\mathbb{C}}^{*}}$; more concretely, with the integers:
$u_{i}^{\pm}:=\dim{\mathcal{U}}_{i}^{\pm}\cap{\mathcal{M}}_{y},\mbox{ $y\in Y_{i}$ general}.$
@@
We start by observing that $B_{i}^{-}\cap B_{i+1}^{+}=q(p^{-1}(p({\mathcal{U}}_{i}^{+}))$, hence
@@ 6. Actions with an isolated extremal fixed point
In the next section we will present some consequences of the assumption that an extremal fixed point of the action is isolated. In particular we will show that the Picard number of every fiber of $q$ is equal to the relative Picard number of $q$ and it is is at most two. We will then divide the proof in two cases, according to its possible values.
@@
Let $B^{-}(Y_{1})$ be the closure of $X^{-}(Y_{1})$; the fact that $\nu^{+}(Y_{1})=1$ (see [25, Lemma 4.5]), implies that $B^{-}(Y_{1})\subset X$ is a divisor. This divisor is locally isomorphic to $N^{-}(Y_{1})$ in a neighbourhood of $Y_{1}$; in particular it is smooth along $Y_{1}$. On the other hand, a bridge $C$ of the family ${\mathcal{M}}$ linking $Y_{1}$ with $y_{0}$ is smooth at its source $y_{1}$ (Lemma 2.3), and meets $B^{-}(Y_{1})$ transversally at $y_{1}$. This implies that $B^{-}(Y_{1})\cdot C=1$, and so $L\cdot C=1$, since $L$ is the ample generator of $\mathop{\rm Pic}\nolimits(X)$. ∎
@@
In the next statement we will identify the variety ${\mathcal{M}}_{y_{0}}$ with a component of fixed points of $X$. We will show that these curves, which are ${\mathbb{C}}^{*}$-invariant, are uniquely determined by their sources, which are the points of $Y_{1}$. This will allow us to describe their universal family as a projective bundle over ${\mathcal{M}}_{y_{0}}$.
@@
${\mathcal{M}}_{y_{0}}\simeq Y_{1},\qquad p^{*}{\mathcal{E}}_{|{\mathcal{M}}_{y_{0}}}\simeq{\mathcal{O}}_{{\mathcal{M}}_{y_{0}}}\oplus q^{*}L_{|{\mathcal{M}}_{y_{0}}}.$
In particular, we have that $L_{|Y_{1}}\simeq D_{|{\mathcal{M}}_{y_{0}}}$.
@@
By Corollary 5.5, $p({\mathcal{U}}^{+}_{0})=p({\mathcal{U}}^{-}_{1})\subset{\mathcal{M}}$. Moreover, the subvarieties ${\mathcal{U}}^{+}_{0},{\mathcal{U}}^{-}_{1}$, of the ${\mathbb{P}}^{1}$-bundle $p^{-1}(p({\mathcal{U}}^{+}_{0}))\to p({\mathcal{U}}^{+}_{0})$ are sections (they correspond, respectively, to the sinks and the sources of the curves parametrized by $p({\mathcal{U}}^{+}_{0})$), therefore they are isomorphic. Since $\nu^{+}(Y_{1})=1$ (see [25, Lemma 4.5]) we have that ${\mathcal{U}}^{-}_{1}\simeq Y_{1}$. On the other hand, by Proposition 5.2, part (e1), the action on ${\mathcal{U}}_{0}={\mathcal{M}}_{y_{0}}$ is trivial, so we conclude that ${\mathcal{U}}_{0}={\mathcal{U}}_{0}^{+}\simeq{\mathcal{U}}_{1}^{-}\simeq Y_{1}$.
Set ${\mathcal{E}}_{0}:=p^{*}{\mathcal{E}}_{|{\mathcal{M}}_{y_{0}}}$. The image of ${\mathcal{M}}_{y_{0}}$ into ${\mathcal{M}}$ via $p$ is the subfamily of ${\mathbb{C}}^{*}$-invariant rational curves with sink $y_{0}$ and source in $Y_{1}$, and ${\mathbb{P}}({\mathcal{E}}_{0})={\mathcal{U}}\times_{{\mathcal{M}}}{\mathcal{M}}_{y_{0}}$ is the universal family of these curves.
There are two disjoint sections, $s_{0}:{\mathcal{M}}_{y_{0}}\to{\mathcal{U}}_{0}^{+}\subset{\mathbb{P}}({\mathcal{E}}_{0})$, $s_{1}:{\mathcal{M}}_{y_{0}}\to{\mathcal{U}}_{1}^{-}\subset{\mathbb{P}}({\mathcal{E}}_{0})$, corresponding to the extremal fixed points of the curves, which map via $q$ to the point $y_{0}$ and to $Y_{1}$, respectively. In particular, ${\mathcal{E}}_{0}$ splits as
${\mathcal{E}}_{0}\simeq s_{0}^{*}q^{*}L_{|{\mathcal{M}}_{y_{0}}}\oplus s_{1}^{*}q^{*}L_{|{\mathcal{M}}_{y_{0}}}.$
Since, via the map $q$, ${\mathcal{U}}_{0}^{+}$ is contracted to $y_{0}$ and ${\mathcal{U}}_{1}^{-}$ maps isomorphically to $Y_{1}$, it follows that ${\mathcal{E}}_{0}\simeq{\mathcal{O}}_{{\mathcal{M}}_{y_{0}}}\oplus q^{*}L_{|{\mathcal{M}}_{y_{0}}}$, from which we finally get that $\mathop{\rm det}\nolimits({\mathcal{E}}_{0})\simeq q^{*}L_{|{\mathcal{M}}_{y_{0}}}\simeq L_{|Y_{1}}$. ∎
@@
Let $f:{\mathbb{P}}^{1}\to C$ be the normalization of $C$. We consider the family of closures of orbits linking the curve $C$ with $y_{0}$, which, as in Lemma 6.4, has universal family ${\mathbb{P}}(f^{*}(p^{*}{\mathcal{E}})_{|{\mathcal{M}}_{y_{0}}}))\simeq{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$. Let us denote by $q^{\prime}:{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1))\to X$ its evaluation map. It contracts the section corresponding to ${\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)\to{\mathcal{O}}_{{\mathbb{P}}^{1}}$ to the point $y_{0}$, and sends the section corresponding to ${\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)\to{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)$ onto the curve $C$. In particular the Stein factorization of $q^{\prime}$ is the contraction of ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$ to ${\mathbb{P}}^{2}$. Then $C$ is algebraically equivalent to the image of any line in ${\mathbb{P}}^{2}$ passing by $y_{0}$, that is to an element of ${\mathcal{M}}$. Since ${\mathcal{M}}$ is a smooth irreducible component of $\mathop{\rm RatCurves}\nolimits^{n}(X)$, it follows that also $C$ is a curve of the family ${\mathcal{M}}$. ∎
@@
The line bundle $D$ has positive degree on rational curves contracted by $q$. Moreover, if $y\in X^{{\mathbb{C}}^{*}}\setminus(Y_{0}\cup Y_{r})$ is an inner fixed point, then the degree of $D$ on the closure of a general orbit of the induced action on ${\mathcal{M}}_{y}$ is two.
@@
Let $\Gamma$ be a rational curve contracted by $q$, denote by $\nu:{\mathbb{P}}^{1}\to p(\Gamma)$ the normalization of $p(\Gamma)$ and consider the fiber square
@@
For the last claim, we notice that, if $y\in X^{{\mathbb{C}}^{*}}\setminus(Y_{0}\cup Y_{r})$ and $\Gamma$ is the closure of a general orbit of the induced action on ${\mathcal{M}}_{y}$, then $e=2$ by Proposition 4.6. ∎
@@
Assume that $\Gamma$ degenerates as $\Gamma_{1}+\Gamma_{2}$. Then, from $(D+q^{*}L)\cdot\Gamma=1$ we get that, for some $i=1,2$, we have $(D+q^{*}L)\cdot\Gamma_{i}\geq 0$, which is impossible, since $q^{*}L$ is nef and $D$ has positive degree on rational curves. ∎
We can now use the above results to bound the Picard number of ${\mathcal{M}}_{y}$ for every point $y$ in $Y_{1}$.
@@
If $y$ is a point of $Y_{1}$ then the action on ${\mathcal{M}}_{y}$ has isolated sink. In particular, ${\mathcal{M}}_{y}$ is rationally connected. Moreover $\rho({\mathcal{M}}_{y})\leq 2$.

### v2

@@ 5.1. The orbit graph
Recalling that a fixed point of ${\mathcal{M}}$ parametrizes either a fixed curve or a bridge, we see that the fixed point components of the action on ${\mathcal{M}}$ are images via $p$ of the varieties ${\mathcal{U}}_{j}^{i}$, $(i=-,0,+)$. We then set:
@@
Let ${\mathcal{N}}$ be an irreducible component of ${\mathcal{M}}_{j}^{0}$. By Remark 2.9, ${\mathcal{N}}$ is an irreducible component of the normalization $\mathop{\rm RatCurves}\nolimits^{n}(Y_{j})$ of $\mathop{\rm RatCurves}\nolimits(Y_{j})$, hence it is a family of rational curves in $Y_{j}$, and it is obviously unsplit. Moreover, since $T_{Y_{j}}$ is a direct summand of ${T_{X}}_{|Y_{j}}$ (see Section 2.1), it follows that $\nu^{*}T_{Y_{j}}$ is nef, for the normalization $\nu$ of any curve of the family ${\mathcal{N}}$, and we conclude that ${\mathcal{N}}$ is beautiful by Remark 2.9. ∎
@@
The proof of this statement will be obtained by combining Theorem 5.3 together with Lemmata 5.8, 5.9, 5.10 below.
@@ 5.2. Geometry of the BB-cells
In this section we will discuss some special properties of the plus and minus BB-decompositions in the assumptions of Setup 5.1. We start by showing that in this situation such decompositions are stratifications, i.e., that the closure of a BB-cell is a union of BB-cells:
@@
It is clearly enough to prove only the first part of the statement, and to prove it only for $j=i+1$. In order to do so, we will show that a general orbit with sink in $Y_{i+1}$ is contained in $B_{i}^{-}$. Note that Theorem 5.7 implies that there exists an orbit with sink in $Y_{i+1}$ and source in $Y_{r}$, from which we may conclude that $B_{i+1}^{-}$ intersects the open subset $X^{+}(Y_{r})$, i.e., that the general orbit with sink in $Y_{i+1}$ has source in $Y_{r}$. Let $\Gamma$ be the closure of such an orbit, and set ${\mathcal{U}}_{\Gamma}:=q^{-1}(\Gamma)$.
@@
Consider the ruled surface $S_{C}$ as in Construction 1. The limit curve $C_{-}$ is a bridge connecting $Y_{i}$ and $Y_{i+1}$, while the limit curve $C_{+}$ is either a curve contained in $Y_{r}$, or a bridge connecting $Y_{r-1}$ and $Y_{r}$ (if $Y_{r}$ is an isolated point). In all cases, the sink of the action on $S_{C}$ is a point mapped in $X$ to a point of $Y_{i}$, and the source is mapped to $Y_{r}$. Hence, a general orbit in $S_{C}$ is mapped to an orbit contained in $B^{-}(Y_{i})$, and therefore $q(S_{C})\subset B^{-}(Y_{i})$; since $\Gamma=q(\Gamma^{\prime})\subset q(S_{C})$, the proof is finished. ∎
@@
In the first case $\Gamma^{\prime}$ connects the source of $\Gamma_{-}$ with the source of $\Gamma_{+}$ and we are in case 0b) of Proposition 4.6; in the second $\Gamma^{\prime}$ connects the source of $\Gamma_{-}$ with a point of $\Gamma_{+}$, which is fixed, and we are in case 1) of Proposition 4.6. In both cases the cycle $\Gamma_{-}+\Gamma^{\prime}$ is a limit of closure of general orbits in $S_{\Gamma}$. These orbits are mapped to orbits with sink $y\in Y_{i}$ and source belonging to $Y_{r}$. ∎
@@
For every $i=0,\dots,r-1$ the intersection $B_{i}^{-}\cap B_{i+1}^{+}$ is proper, i.e., $\mathop{\rm codim}\nolimits(B_{i}^{-}\cap B_{i+1}^{+})=\mathop{\rm codim}\nolimits B_{i}^{-}+\mathop{\rm codim}\nolimits B_{i+1}^{+}=\nu^{+}_{i}+\nu^{-}_{i+1}$.
@@
Let us consider again the Chow quotient $\mathcal{C}\!X$ of $X$ by the ${\mathbb{C}}^{*}$-action (cf. [24, Section 4]), which is an $(n-1)$-dimensional projective variety.
Set $B:=B_{i}^{-}\cap B_{i+1}^{+}$. Since, by Theorem 5.3, there exists a bridge between $Y_{i}$ and $Y_{i+1}$, then $B$ intersects the open sets $X^{-}(Y_{i})\subset B_{i}^{-}$, $X^{+}(Y_{i+1})\subset B_{i+1}^{+}$, and we may claim that the ${\mathbb{C}}^{*}$-orbit of a general point in $B$ has its extremal fixed point components in $Y_{i},Y_{i+1}$. We then conclude that $B$ is the locus of the set of bridges between $Y_{i}$ and $Y_{i+1}$, that we denoted by $I$; note that $I={\mathcal{U}}_{i}^{+}={\mathcal{U}}_{i+1}^{-}$. In particular $B$ is irreducible and $\dim B=\dim I+1$.
We denote by $e_{i}:I\to Y_{i}$ (resp. $e_{i+1}:I\to Y_{i+1}$) the map assigning to every element of $I$ its sink (resp. its source). Let us denote by $V_{i}^{+}\subset{\mathbb{P}}(N^{+}(Y_{i}))$, (resp. $V_{i+1}^{-}\subset{\mathbb{P}}(N^{-}(Y_{i+1}))$) the open set parametrizing orbits linking $Y_{0}$ and $Y_{i}$ (resp. $Y_{i+1}$ and $Y_{r}$), which is non-empty by Theorem 5.7. Consider the fiber product:
@@
which parametrizes all the connected ${\mathbb{C}}^{*}$-invariant $1$-cycles of the form $C^{\prime}+C+C^{\prime\prime}$, where $C^{\prime}$ is the closure of an orbit linking $Y_{0}$, $Y_{i}$, $C\in I$ and $C^{\prime\prime}$ is the closure of an orbit linking $Y_{i+1}$, $Y_{r}$. The variety $Z$ is smooth, irreducible, of dimension:
@@
On the other hand, we may consider in $\mathcal{C}\!X$ the subvariety parametrizing connected ${\mathbb{C}}^{*}$-invariant cycles of the form $C_{1}+C_{2}$ with $C_{1}$ linking $Y_{0}$ and $Y_{i}$, $C_{2}$ linking $Y_{i}$ and $Y_{r}$. It is isomorphic to an open subset of ${\mathbb{P}}(N^{-}(Y_{i}))\times_{Y_{i}}{\mathbb{P}}(N^{+}(Y_{i}))$. We denote its closure by $E_{i}\subset\mathcal{C}\!X$, and define analogously $E_{i+1}$. By construction $E_{i}$, $E_{i+1}$ are irreducible divisors in $\mathcal{C}\!X$.
@@
In order to prove that $Z\subset E_{i}\cap E_{i+1}$ we have to show that every cycle of $Z$ belongs to $E_{i}\cap E_{i+1}$. Such cycles are of the form $C^{\prime}+C+C^{\prime\prime}\in Z$, where $C^{\prime},C,C^{\prime\prime}$ are ${\mathbb{C}}^{*}$-invariant irreducible curves linking, respectively, $Y_{0}$ to $Y_{i}$, $Y_{i}$ to $Y_{i+1}$, and $Y_{i+1}$ to $Y_{r}$. It is then enough to notice that, by Lemma 5.12, $C+C^{\prime\prime}$ deforms algebraically to a ${\mathbb{C}}^{*}$-invariant curve linking $Y_{0}$ to $Y_{i+1}$ and, analogously, that $C^{\prime}+C$ deforms algebraically to a ${\mathbb{C}}^{*}$-invariant curve linking $Y_{i}$ to $Y_{r}$. ∎
An important outcome of the previous statement, that we will use later on, is that it allows us to relate the ranks $\nu^{\pm}_{i}$ to the dimensions of the extremal fixed point components of the action on the varieties ${\mathcal{U}}_{y}$, $y\in X^{{\mathbb{C}}^{*}}$; more concretely, with the integers:
$u_{i}^{\pm}:=\dim{\mathcal{U}}_{i}^{\pm}\cap{\mathcal{U}}_{y},\mbox{ $y\in Y_{i}$ general}.$
@@
We start by observing that $B_{i}^{-}\cap B_{i+1}^{+}=q(p^{-1}(p({\mathcal{U}}_{i}^{+})))$, hence
@@ 6. Actions with an isolated extremal fixed point
In the next section we will present some consequences of the assumption that an extremal fixed point of the action is isolated. We will then divide the proof in two cases, according to the possible values of the relative Picard number $\rho({\mathcal{U}}/X)$.
@@
Let $B^{-}(Y_{1})$ be the closure of $X^{-}(Y_{1})$; the fact that $\nu^{+}(Y_{1})=1$ (see [23, Lemma 4.5]), implies that $B^{-}(Y_{1})\subset X$ is a divisor. This divisor is locally isomorphic to $N^{-}(Y_{1})$ in a neighbourhood of $Y_{1}$; in particular it is smooth along $Y_{1}$. On the other hand, a bridge $C$ of the family ${\mathcal{M}}$ linking $Y_{1}$ with $y_{0}$ is smooth at its source $y_{1}$ (Lemma 2.3), and meets $B^{-}(Y_{1})$ transversally at $y_{1}$. This implies that $B^{-}(Y_{1})\cdot C=1$, and so $L\cdot C=1$, since $L$ is the ample generator of $\mathop{\rm Pic}\nolimits(X)$. ∎
@@
In the next statement we will identify the variety ${\mathcal{U}}_{y_{0}}$ with a fixed point component of $X$. In fact, we will show that the curves of the family ${\mathcal{M}}_{y_{0}}$ are uniquely determined by their sources, which are the points of $Y_{1}$; we also know that the curves of ${\mathcal{M}}_{y_{0}}$ are smooth at $y_{0}$, then the universal family over ${\mathcal{M}}_{y_{0}}$ contains two sections: one is ${\mathcal{U}}_{y_{0}}$ and the other is isomorphic to $Y_{1}$.
@@
${\mathcal{U}}_{y_{0}}\simeq Y_{1},\qquad p^{*}{\mathcal{E}}_{|{\mathcal{U}}_{y_{0}}}\simeq{\mathcal{O}}_{{\mathcal{U}}_{y_{0}}}\oplus q^{*}L_{|{\mathcal{U}}_{y_{0}}}.$
In particular, we have that $L_{|Y_{1}}\simeq D_{|{\mathcal{U}}_{y_{0}}}$.
@@
By Corollary 5.5, $p({\mathcal{U}}^{+}_{0})=p({\mathcal{U}}^{-}_{1})\subset{\mathcal{M}}$. Moreover, the subvarieties ${\mathcal{U}}^{+}_{0},{\mathcal{U}}^{-}_{1}$, of the ${\mathbb{P}}^{1}$-bundle $p^{-1}(p({\mathcal{U}}^{+}_{0}))\to p({\mathcal{U}}^{+}_{0})$ are sections (they correspond, respectively, to the sinks and the sources of the curves parametrized by $p({\mathcal{U}}^{+}_{0})$), therefore they are isomorphic. Since $\nu^{+}(Y_{1})=1$ (see [23, Lemma 4.5]) we have that ${\mathcal{U}}^{-}_{1}\simeq Y_{1}$. On the other hand, by Proposition 5.2, part (e1), the action on ${\mathcal{U}}_{0}={\mathcal{U}}_{y_{0}}$ is trivial, so we conclude that ${\mathcal{U}}_{0}={\mathcal{U}}_{0}^{+}\simeq{\mathcal{U}}_{1}^{-}\simeq Y_{1}$.
Set ${\mathcal{E}}_{0}:=p^{*}{\mathcal{E}}_{|{\mathcal{U}}_{y_{0}}}$. The image ${\mathcal{M}}_{y_{0}}$ of ${\mathcal{U}}_{y_{0}}$ into ${\mathcal{M}}$ via $p$ is the subfamily of ${\mathbb{C}}^{*}$-invariant rational curves with sink $y_{0}$ and source in $Y_{1}$, and ${\mathbb{P}}({\mathcal{E}}_{0})={\mathcal{U}}\times_{{\mathcal{M}}}{\mathcal{U}}_{y_{0}}$ is the universal family of these curves.
There are two disjoint sections, $s_{0}:{\mathcal{U}}_{y_{0}}\to{\mathcal{U}}_{0}^{+}\subset{\mathbb{P}}({\mathcal{E}}_{0})$, $s_{1}:{\mathcal{U}}_{y_{0}}\to{\mathcal{U}}_{1}^{-}\subset{\mathbb{P}}({\mathcal{E}}_{0})$, corresponding to the extremal fixed points of the curves, which map via $q$ to the point $y_{0}$ and to $Y_{1}$, respectively. In particular, ${\mathcal{E}}_{0}$ splits as
${\mathcal{E}}_{0}\simeq s_{0}^{*}q^{*}L_{|{\mathcal{U}}_{y_{0}}}\oplus s_{1}^{*}q^{*}L_{|{\mathcal{U}}_{y_{0}}}.$
Since, via the map $q$, ${\mathcal{U}}_{0}^{+}$ is contracted to $y_{0}$ and ${\mathcal{U}}_{1}^{-}$ maps isomorphically to $Y_{1}$, it follows that ${\mathcal{E}}_{0}\simeq{\mathcal{O}}_{{\mathcal{U}}_{y_{0}}}\oplus q^{*}L_{|{\mathcal{U}}_{y_{0}}}$, from which we finally get that $\mathop{\rm det}\nolimits({\mathcal{E}}_{0})\simeq q^{*}L_{|{\mathcal{U}}_{y_{0}}}\simeq L_{|Y_{1}}$. ∎
@@
Let $f:{\mathbb{P}}^{1}\to C$ be the normalization of $C$. We consider the family of closures of orbits linking the curve $C$ with $y_{0}$, which, as in Lemma 6.4, has universal family ${\mathbb{P}}(f^{*}(p^{*}{\mathcal{E}})_{|{\mathcal{U}}_{y_{0}}}))\simeq{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$. Let us denote by $q^{\prime}:{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1))\to X$ its evaluation map. It contracts the section corresponding to ${\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)\to{\mathcal{O}}_{{\mathbb{P}}^{1}}$ to the point $y_{0}$, and sends the section corresponding to ${\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)\to{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)$ onto the curve $C$. In particular the Stein factorization of $q^{\prime}$ is the contraction of ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$ to ${\mathbb{P}}^{2}$. Then $C$ is algebraically equivalent to the image of a line in ${\mathbb{P}}^{2}$ passing by $y_{0}$, that is to an element of ${\mathcal{M}}$. Since ${\mathcal{M}}$ is a smooth irreducible component of $\mathop{\rm RatCurves}\nolimits^{n}(X)$, then $C$ is a curve of the family ${\mathcal{M}}$, as well. ∎
@@
The line bundle $D$ has positive degree on rational curves contracted by $q$. Moreover, if $y\in X^{{\mathbb{C}}^{*}}\setminus(Y_{0}\cup Y_{r})$ is an inner fixed point, then the degree of $D$ on the closure of a general orbit of the induced action on ${\mathcal{U}}_{y}$ is two.
@@
Let $\Gamma$ be a rational curve contracted by $q$, denote by $\nu:{\mathbb{P}}^{1}\to p(\Gamma)$ the normalization of $p(\Gamma)$ and consider the Cartesian square
@@
For the last claim, we notice that, if $y\in X^{{\mathbb{C}}^{*}}\setminus(Y_{0}\cup Y_{r})$ and $\Gamma$ is the closure of a general orbit of the induced action on ${\mathcal{U}}_{y}$, then $e=2$ by Proposition 4.6. ∎
@@
Assume that $\Gamma$ degenerates as $\Gamma_{1}+\Gamma_{2}$. Then, from $(D+q^{*}L)\cdot\Gamma=1$ we get that, for some $i=1,2$, we have $(D+q^{*}L)\cdot\Gamma_{i}\leq 0$, which is impossible, since $q^{*}L$ is nef and $D$ has positive degree on rational curves. ∎
We can now use the above results to bound the Picard number of ${\mathcal{U}}_{y}$ for every point $y$ in $Y_{1}$.
@@
If $y$ is a point of $Y_{1}$ then the action on ${\mathcal{U}}_{y}$ has isolated sink. In particular, ${\mathcal{U}}_{y}$ is rationally connected. Moreover $\rho({\mathcal{U}}_{y})\leq 2$.